OpenStudy (anonymous):

How do you simplify (6 x 10^3)(5 x 10^-4) ?

OpenStudy (anonymous):

In multiplication you can rearrange the terms you are allowed to do that.. So: \[(6 \times 10^{3})(5 \times 10^{-4}) \implies (6 \times 5) \cdot (10^3 \times 10^{-4})\]

OpenStudy (anonymous):

Tell me what do you get: \(6 \times 5\) = ??

OpenStudy (anonymous):

@hongkongfuiz1234 I am waiting for your reply...

OpenStudy (anonymous):

Are you interested in replying or not @hongkongfuiz1234

OpenStudy (anonymous):

Sorry, I left for school. Thank you for helping though! :D